OpenStudy (cassieforlife5):

find where the slope of the curve is defined? x= cos y I've already found the implicit differentiation -1/siny but I don't know what to do from here

OpenStudy (xapproachesinfinity):

x=cosy the slope is \[\frac{dx}{dy}=-\sin y \] this defined for any value of y so why do you need implicit derivative unless you seeking dy/dx

OpenStudy (xapproachesinfinity):

if you are lookinf ro dy/dx you are esentially looking for the derivative of y=arccos(x) \[x=\cos y \Longrightarrow 1=\frac{dy}{dx}(-\sin y) \] which is what you got \[\frac{dy}{dx}=-\frac{1}{\sin y}\] this defined for \[y=0, \pi, 2\pi... \] if we generlize \[y' \text{is defined for the set of values y that satisfy the following:}\] \[\ \{y: y\ne 2n\pi , ~~n\in \mathbb{Z} \}\ \]